Hello

Hope all is well, Happy New Year. I was wondering if you can help me with a math question that i'm having a little trouble with. I'm not sure if i use the 10 or 100 to mutiply it with. I will show you what i have to far hopefully it is right. Thank you so much for your help.
Question:
How many ounces of a 50% antifreeze solution must be mixed with 10 oz of an 80% antifreeze solution to produce a 60% antifreeze solution?
What i have so far:
.50+.80=.60
.50(x)+.80(10)=.60(x+10)
50+( This is where i get mixed up and im not sure my next move. Multipy by 10 or 100.If you can help me i will appreciate it alot thanks again.)

Answer by stanbon(75887)

How many ounces of a 50% antifreeze solution must be mixed with 10 oz of an 80% antifreeze solution to produce a 60% antifreeze solution?
-------
Equation:
anti + anti = anti
0.80*10 + 0.50x = 0.60(10+x)
----
Multiply thru by 100 to get:
80*10 + 50x = 60*10 + 60x
10x = 20*10
x = 20 oz (amt. of 50% anti needed)
=======
Cheers,
Stan H.